The length of a rectangle is twice its width.
If the perimeter of the rectangle is 60 yds , find its area.

Answer by checkley77(12844)

L=2W
2L+2W=P
2(2W)+2W=60
4W+2W=60
6W=60
W=60/6
W=10 ANS. FOR THE WIDTH.
L=2*10=20 ANS. FOR THE LENGTH.
LW=A
20*10=A
200=A ANSWER FOR THE AREA.